Hi, I'm having issues understanding a quadratic equation that was given to me as an example. I do not under stand the second line and how they got to it. Could someone please explain it or show the steps that were taken to get there.
The equation is d^2-7d+9=0. Here are the steps I was given.
d^=7d-9
d^2-7d+9=0
d=(7+_ Square root of 49-4(9)) divided by 2
d=(7+_squareroot of 49-36)/2
d= (7+- squareroot of 13) /2
Now I can see how step 1 to 2 is done that is simple and I can see how it goes from step 4 to the answer but I can not understand how to go from step 2 to step 3.
I have tried it myself. here is what I keep getting.
d^2=7d-9
d^2-7d+9=0
d(d-7)+9=0
d(d-7)=-9
d-7=-9/d
d=-9/d+7
2d=-9+7
2d=-2
d=-1
I know my answer is incorrect, Please Help me

Answer by mananth(16946)

The equation is d^2-7d+9=
add -9 to both sides
d^2-7d+9-9 =-9
we get d^2-7 = -9
Add a third term on the left side such that it becomes a perfect square. The same term you add on the right side
d^2-7d+49/4 = -9 +(49/4)
(d-7/2)^2= -9+(49/4)( take the LCD )
(d-7/2)^2=(-36+49)/4
(d-7/2)^2=13/4
Take the square root on both sides
d-7/2= +/- sqrt(13)/sqrt(4)
d1= 7/2+sqrt(13)/2
d=5.03
OR
d2= 7/2-sqrt(13)/2
1.697